the curves y= 4x^2 + 2x -8 and y= x^3 -x +13 touch eachother at ......?

Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
The condition for the two curves to touch each other is that they should both have a common tangent at a particular (x,y) :
=> meet at (x1,y1).
dy/dx = 8x1 + 2
dy/dx = 3(x1)^2 - 1
=>3(x1)^2 - 1 = 8x1 + 2
=>3(x1)^2 - 8x1 - 3 =0
x1 = 3, -1/3
y1=34
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty